OpenStudy (anonymous):

What is an equation of a parabola with a vertex at the origin and directrix x = 4.75?

OpenStudy (anonymous):

\[x^2=-4py\] with \(p=4.75\)

OpenStudy (anonymous):

since vortex is at the (0,0) so y^2=-4px p=4.75 y^2=-19x

OpenStudy (anonymous):

So what am i susposed to do ?

OpenStudy (anonymous):

@sami-21 is right and i am wrong i got the square wrong

OpenStudy (anonymous):

\[y^2=-19x\]

OpenStudy (anonymous):

So would it be -1/19y^2 ?

OpenStudy (anonymous):

if you solve for \(x\) it would be \[x=-\frac{1}{19}y^2\] yes

OpenStudy (anonymous):

yes thank you. Can you have help with 3 more ?